OpenStudy (anonymous):

marlisa has exactly $61 in onedollar five dollar and ten dollar bills if she has 14bills in all how many of each bills does she have.

OpenStudy (anonymous):

Does she need to have one of each at least? Can we say that she has no 5 dollars bills?

OpenStudy (anonymous):

idk probably anything i wrote the problem word for word from my book

OpenStudy (anonymous):

ok so assume she must 1 of each at least. 1 + 5 + 10 = 16 dollars subtract that from the total 61-16=45 And we have 11 bills left.

OpenStudy (anonymous):

6= $1 5= $5 3= $10

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

6= $1 5= $5 3= $10 That is the answer!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ty you are very good helper i am fan of yours

OpenStudy (anonymous):

Assume then that we have all 5 dollar bills left. 5* 11=55 Wow we can't have 55 we need to have 45 that's way too much. So we need dollar bills too. Ok lets then say we have 5 one dollar bills That leaves us 45-5=40 dollars ok now 11-5= 6 bills left Now we can make a combination of 4 5dollar bills =20 ans we have 2 bills of 10= 20 20+20 =40 1+5=6 1$ 1+4=5 5$ 1+2=3 10$

OpenStudy (anonymous):

i will be back on in a couple days but i appreciate all the help and you are da bom romero.